Ben has more coins than William in 7:5 ratio. Ben gives William 12 of his. William, with those 12, has the same number as Ben.
Determine number of coins each had before Ben gave William the coins.
Not sure how to solve.

Answer by ikleyn(52905)

.

Initially, Ben had 7x coins;  William had 5x coins.


After giving 12 coins from Ben to William


    7x - 12 = 5x + 12


Simplify and find x


    7x - 5x = 12 + 12

       2x   =    24

        x    =   24/2 = 12.


ANSWER.  Initially, Ben had 7*12 = 84 coins;  William had 5*12 = 60 coins.

Solved.